### v1

@@ Abstract
One of the prime problems of computer science and machine learning is to extract information efficiently from large-scale, heterogeneous data. Text data, with its syntax, semantics, and even hidden information content, possesses an exceptional place among the data types in concern. The processing of the text data requires embedding, a method of translating the content of the text to numeric vectors. A correct embedding algorithm is the starting point for obtaining the full information content of the text data. In this work, a new embedding method based on the graph structure of the meaningful sentences is proposed. The design of the algorithm aims to construct an embedding vector that constitutes syntactic and semantic elements as well as the hidden content of the text data. The success of the proposed embedding method is tested in classification problems. Among the wide range of application areas, text classification is the best laboratory for embedding methods; the classification power of the method can be tested using dimensional reduction without any further processing. Furthermore, the method can be compared with different embedding algorithms and machine learning methods. The proposed method is tested with real-world data sets and eight well-known and successful embedding algorithms. The proposed embedding method shows significantly better classification for binary and multiclass datasets compared to well-known algorithms.
@@
GuideWalk - Heterogeneous Data Fusion for Enhanced Learning - A Multiclass Document Classification Case
@@ 3.1 Universal weighted graph of word
Figure 2: Degree distribution of words in the generated word-graph.
@@ 4 Results
where $TP_{s}$, $TN_{s}$, $FP_{s}$, and $FN_{s}$ represent the number of true positives, true negatives, false positives and false negatives, respectively, for category $s$ within the set of labels $S$.
@@ 4.1 The Datasets
To obtain node feature vectors, the proposed algorithm (GTPM) applies the following procedure to each dataset: After the cleaning, the distinct words become nods of the graph. Connections and the connection weights are calculated using the information embedded in the original text. Random walks were performed on the weighted graph to calculate the transition probability matrix (TPM) [23]. The obtained transition probability matrix is converted to distinct feature vectors for each node, representing the local topology and connectivity of the starting node.
@@ 4.2 Baseline Models
  • TF-IDF + LR: In the process of feature extraction, TF-IDF is employed to transform the textual documents into feature representations. These TF-IDF-based features are subsequently employed as inputs for the Logistic Regression (LR) classifier.
@@
  • Bi-LSTM: Bidirectional Long Short-Term Memory (Bi-LSTM) is a type of recurrent neural network architecture designed to address the vanishing gradient problem common in standard RNNs. Bi-LSTMs process sequence data in both forward and backward directions, enabling the capture of long-term dependencies. For this study, a Bi-LSTM model was implemented utilizing pre-trained BERT word embeddings as input features. The final hidden state of the Bi-LSTM was used for text classification.
@@
  • TextGCN: TextGCN [35] models the text corpus as a single interconnected graph with words and documents as nodes and document-word and word-word co-occurrence as edges, then utilizes graph convolutional networks to perform node classification on this graph representation for text classification.
  • TextING: The TextING model constructs separate graphs for each document and utilizes a gated graph neural network to learn text-level word interactions, aiming to improve its ability to preserve both local and global contextual information in an inductive learning fashion [36].
@@
  • VGCN-BERT: A graph convolutional network extracts features from a vocabulary graph to enhance word embeddings, which are input to the self-attention encoder in BERT for classification [38].
@@ 4.5.1 Optimal Parameter Selection
The observed behavior is that the model’s effectiveness improves with the increasing walk length up to a point, and it decreases after a peak. The model achieves optimal performance when the walk length is around 15 steps. The optimal number of random walks started from each node (word in a document) depends on the dataset’s average text length. For datasets with longer average text lengths, such as Reuters and 20NG, initiating a single random walk from each node yielded the best performance. Conversely, datasets with shorter average text lengths, such as MR and CoLA, require more than one random walk per word for better results. In the comparative studies of the success of the embedding, the length of walks is set to $m=15$ for the GTPM. Similarly, the performances of all other embedding algorithms used for comparisons were measured using the optimal parameter values obtained from the literature.

### v2

@@ Abstract
One of the prime problems of computer science and machine learning is to extract information efficiently from large-scale, heterogeneous data. Text data, with its syntax, semantics, and even hidden information content, possesses an exceptional place among the data types in concern. The processing of the text data requires embedding, a method of translating the content of the text to numeric vectors. A correct embedding algorithm is the starting point for obtaining the full information content of the text data. In this work, a new text embedding approach, namely the Guided Transition Probability Matrix (GTPM) model is proposed. The model uses the graph structure of sentences to capture different types of information from text data, such as syntactic, semantic, and hidden content. Using random walks on a weighted word graph, GTPM calculates transition probabilities to derive text embedding vectors. The proposed method is tested with real-world data sets and eight well-known and successful embedding algorithms. GTPM shows significantly better classification performance for binary and multi-class datasets than well-known algorithms. Additionally, the proposed method demonstrates superior robustness, maintaining performance with limited (only $10\%$) training data, showing an $8\%$ decline compared to $15-20\%$ for baseline methods.
@@
GuideWalk: A Novel Graph - Based Word Embedding for Enhanced Text Classification
@@ 3.1 Universal weighted graph of word
Figure 2: Degree distribution of words in the generated word-graph using the Reuters dataset.
@@ 4 Results
where $S$ represents the number of classes in a multi-class classification problem. $P_{s}$ and $R_{s}$ represent precision of a specific class $s$ and recall of a specific class $s$. $TP_{s}$, $TN_{s}$, $FP_{s}$, and $FN_{s}$ represent the number of true positives, true negatives, false positives and false negatives, respectively, for category $s$ within the set of classes $S$.
@@ 4.1 The Datasets
To obtain node feature vectors, the proposed algorithm (GTPM) applies the following procedure to each dataset: After the cleaning, the distinct words become nodes of the graph. Connections and the connection weights are calculated using the information embedded in the original text. Random walks were performed on the weighted graph to calculate the transition probability matrix (TPM) [23]. The obtained transition probability matrix is converted to distinct feature vectors for each node, representing the local topology and connectivity of the starting node.
@@ 4.2 Baseline Models
  • TF-IDF + LR: In the process of feature extraction, Term Frequency-Inverse Document Frequency (TF-IDF) is employed to transform the textual documents into feature representations. These TF-IDF-based features are subsequently employed as inputs for the Logistic Regression (LR) classifier.
@@
  • Bi-LSTM: Bidirectional Long Short-Term Memory (Bi-LSTM) is a type of recurrent neural network architecture designed to address the vanishing gradient problem common in standard RNNs. Bi-LSTMs process sequence data in both forward and backward directions, enabling the capture of long-term dependencies. For this study, a Bi-LSTM model was implemented utilizing pre-trained Bidirectional Encoder Representations from Transformers (BERT) word embeddings as input features. The final hidden state of the Bi-LSTM was used for text classification.
@@
  • TextGCN: Text Graph Convolutional Network (TextGCN) [35] models the text corpus as a single interconnected graph with words and documents as nodes and document-word and word-word co-occurrence as edges, then utilizes graph convolutional networks to perform node classification on this graph representation for text classification.
  • TextING: Text classification method for INductive word representations via Graph neural networks (TextING) constructs separate graphs for each document and utilizes a gated graph neural network to learn text-level word interactions, aiming to improve its ability to preserve both local and global contextual information in an inductive learning fashion [36].
@@
  • VGCN-BERT: Vocabulary Graph Convolutional Network-BERT (VGCN-BERT) model uses a simple graph convolutional network to extract features from a vocabulary graph to enhance word embeddings, which are input to the self-attention encoder in BERT for classification [38].
@@ 4.5.1 Optimal Parameter Selection
The observed behavior is that the model’s effectiveness improves with the increasing walk length up to a point, and it decreases after a peak. The model achieves optimal performance when the walk length is around 15 steps. The optimal number of random walks started from each node (word in a document) depends on the dataset’s average sentence length. For datasets with longer average sentence lengths, such as Reuters and 20NG, initiating a single random walk from each node yielded the best performance. Conversely, datasets with shorter average sentence lengths, such as MR and CoLA, require more than one random walk per word for better results. In the comparative studies of the success of the embedding, the length of walks is set to $m=15$ for the GTPM. Similarly, the performances of all other embedding algorithms used for comparisons were measured using the optimal parameter values obtained from the literature.